indicated operation: y-2 over 3 = 2y+1 over 5 _
Answer by CubeyThePenguin(3113)

(y-2)/3 = (2y+1)/5
5(y - 2) = 3(2y + 1)
5y - 10 = 6y + 3
-13 = y

y = -13